You carry out a serial dilution of a bacterial culture, using dilutions of ⅕, ⅕, 1/100, and 1/10. Then you plate 100 µl of the final dilution and the next day observe 200 colonies on your plate. What was the concentration of the original culture? (5 pts)
Answer #1

colony count = 200

dilution = 1/5 x 1/5 x 1/100 x 1/10 = 1/25000

100µl of final dilution (1/25000) plated

100µl = 1/10ml

Therefore,

1/10 ml of 1/25000 dilution gives 200 CFU

0.1ml x 0.00004 dilution gives 200 CFU

            0.000004 ml gives 200 CFU

1ml gives 200/0.000004 CFU

= 5 x 107 CFU

Concentration of original culture is 5 x 107 CFU/ml